OpenStudy (anonymous):

1) Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = - 1/2 x + 4 1. y = 2x + 8 2. y = 2x 3. y = -2x - 8 4. y = -2x 2) Choose the equation of the line passing through the point (-2, 6) and parallel to y = 1/2 x - 8 1. y = 1/2 x + 7 2. y = 1/2 x + 5 3. y = 2x + 7 4. y = 2x + 5 3) Choose the equation of the horizontal line that passes through the point (-8, -7) 1. y = -8 2. y = -7 3. x = -8 4. x = -7

OpenStudy (anonymous):

1) perpendicular= opp. reciprocal (flip fraction) of slope slope=2 y=2x+b (plug in x,y into this) x,y is (-2,4) in this case 4=2(-2)+b b= 8 y=2x+8 2)do the same thing as above EXCEPT keep the slope the same because parallel= same slope as the equation so, y=1/2x+b plug in x,y 6=1/2(-2)+b b= 7 y=1/2x+7 i think you are missing 1/2x in choice 1. 3) when y= (any value) it's an horizontal when x= (any value) it's a vertical line (-8,-7) so (x,y) for a HORIZONTAL line, it would be y= -7

OpenStudy (anonymous):

Yeah, on the first two options in #2 I forgot to put the 1/2 in front of the x. Anyways, thank you so much for helping me with those problems! I'll give you a medal too. :)

OpenStudy (anonymous):

ty, have a nice day :D

OpenStudy (anonymous):

Welcome, you too!